# Improving exponential-family random graph models for bipartite networks

### v1

@@ 6 Simulation experiments
Figure 4: Effect of varying the BipartiteAltKCyclesA parameter (left) and BipartiteFourCyclesNodePowerA parameter (right) on the Edge statistic (top), the statistic corresponding to the parameter itself (middle) and the FourCycles statistic (bottom). Each graph plots the mean value of the statistic (over 100 simulated networks) for three different values of the $\lambda$ or $\alpha$ parameter for BipartiteAltKCyclesA and BipartiteFourCyclesNodePowerA, respectively.
6.1 Interpretation of the new parameters
Interpretation of the BipartiteFourCyclesNodePowerA and BipartiteFourCyclesNodePowerB parameters is rather more complicated and is illustrated in Fig. 5 and Fig. 6. Note that in a bipartite network, any four-cycle must contain two nodes in node set $A$ and two nodes and in node set $B$. So how can we get more four-cycles in one mode than the other? The answer is that the four-cycle counts for the two must be equal, but the weighted node-oriented four-cycle counts (8) and (9) can differ. As discussed in Section 5, the statistic BipartiteFourCyclesNodePowerA (8) is maximised by having four-cycles involving distinct pairs of nodes in node set $A$ (rather than many four-cycles involving the same node in node set $A$). If the BipartiteFourCyclesNodePower parameter for $A$ is positive and for $B$ is zero (or negative) then we tend to get more mode $A$ nodes involved in four-cycles, with the same mode $B$ nodes participating in many four-cycles, since the statistic is higher by having different nodes in the four-cycles, than for having the same node involved in many four-cycles. In the examples illustrated in Fig. 6, this results in the mode $A$ nodes being part of a denser core with lots of four-cycles with a smaller number of $B$ nodes, resulting in isolated $B$ nodes. And vice versa for $A$ zero (or negative) and $B$ positive (“zero.pos” and “neg.pos”; these are perhaps clearer as there are more $A$ nodes than $B$ nodes in the network). Particularly in the “neg.pos” case, we can see a core of mode $B$ nodes connected to a small number of central mode $A$ nodes in four-cycles (as well as others not involved in four-cycles) and many mode $A$ isolates. If BipartiteFourCyclesNodePower for both $A$ and $B$ are positive then there are even more four-cycles, but they are more evenly distributed between the $A$ and $B$ nodes.
Figure 5: Effect of different combinations of negative, zero, and positive values of the BipartiteFourCyclesNodePowerA and BipartiteFourCyclesNodePowerB parameters on (top to bottom) the BipartiteFourCyclesNodePowerA statistic, the BipartiteFourCyclesNodePowerB statistic, and the FourCycles statistic. Box plots show the statistics of 100 simulated networks. The simulated bipartite networks have 100 nodes in node set A and 50 nodes in node set B and are simulated with common parameters $(\mathrm{Edge},\mathrm{BipartiteAltStarsA}~{}[\lambda=2],\mathrm{BipartiteAltStarsB}~{}[\lambda=2])=(-6.0,-0.4,1.0)$. For $\mathrm{BipartiteFourCyclesNodePowerA}~{}[\alpha=1/5]$ and $\mathrm{BipartiteFourCyclesNodePowerB}~{}[\alpha=1/5]$ the negative (“neg”) parameter value is $-1.5$ and the positive (“pos”) parameter value is $6.5$.
7 Empirical applications
In this section we demonstrate the application of the new statistics in ERGM models of a selection of empirical two-mode networks ranging in size from hundreds of nodes to hundreds of thousands of nodes. Network summary statistics of the empirical networks are shown in Table 4 (four-cycle counts are shown in Table E1). The largest of these networks is larger than any of the networks in the literature survey (Table LABEL:tab:literature).
Table 4: Bipartite graph summary statistics. “Bipartite c.c” is the Robins-Alexander bipartite clustering coefficient [9].
Network $\mathrm{N_{A}}$ $\mathrm{N_{B}}$ Mean degree Bipartite
A B c.c.
Inouye-Pyke pollinators 91 42 3.09 6.69 0.20856
St Louis crime 870 557 1.71 2.67 0.05041
Norwegian director interlock 1 495 367 1.23 5.00 0.09651
Robertson pollinators 1 428 456 10.68 33.45 0.10876
Australian director interlock 9 971 2 087 1.35 6.45 0.08438
Scientific collaborations 16 726 22 015 3.50 2.66 0.22822
International director interlock 321 869 34 769 1.17 10.84 0.12617
Three of the empirical network examples are director interlock networks [88], where nodes represent directors and companies, with an edge indicating that a director sits on a company board. Two of the empirical network examples are plant-pollinator webs, where edges represent an interaction between a plant and a pollinator, one network represents participation of actors in crime events, and one network represents scientific collaboration (co-authorship of scientific papers).
All of the ERGM models in this section were estimated with EstimNetDirected [71] using the simplified EE algorithm [28] with the improved fixed density (IFD) sampler [26] with 64 parallel estimation runs. Only models in which all runs converge and for which the model is non-degenerate (the observed statistics for effects in the model are with the 95% region of statistics simulated from the model) are presented. Simulations from the estimated models, used for the degeneracy check and goodness-of-fit testing, were done with the SimulateERGM program from the EstimNetDirected software package, with the tie/no-tie (TNT) sampler [87], with the exception of the Inouye-Pyke pollinator web, in which the IFD sampler was used.
7.1 Inouye-Pyke pollinator web
The Inouye-Pyke pollinator web [89] data was downloaded from the Interaction Web Database (IWDB) at https://www.nceas.ucsb.edu/interactionweb/html/inouye_1988.html (accessed 5 June 2018). (Note that the IWDB has since relocated and this data is now available from http://www.ecologia.ib.usp.br/iwdb/html/inouye_1988.html.) In this network, the node set $A$ represents pollinators and node set $B$ represents plants, and an edge represents an interaction between a pollinator and a plant.
Table 5 shows two ERGM models of Inouye-Pyke pollinator web. Model 1 includes the BipartiteAltKCyclesB term (models with BipartiteAltKCyclesA or FourCycles did not converge). Model 2 includes the new BipartiteFourCyclesNodePowerA and BipartiteFourCyclesNodePowerB terms. The former is found to be negative (but not statistically significant) while the latter is positive and statistically significant. This indicates a statistically significant tendency towards having multiple four-cycles involving distinct pairs of plants (node set $B$), and together with the negative or zero (not significant) BipartiteFourCyclesNodePowerA parameter, it models a network with pollinators (node set $A$) participating in multiple four-cycles. This is similar to the “zero.pos” or “neg.pos” case in Fig. 5 and Fig. 6, not coincidentally, since that simulation was based on this model.
Table 5: ERGM parameter estimates for the Inouye-Pyke pollinator web with 95% confidence intervals. Estimates in bold are statistically significant at this level. “Bipartite” in effect names is abbreviated to “Bp”.
Effect Model 1 Model 2
Edge $\mathbf{\underset{(-6.366,-5.598)}{-5.982}}$ $\mathbf{\underset{(-6.470,-5.989)}{-6.229}}$
BpAltStarsA [$\lambda=2$] $\underset{(-0.552,2.397)}{0.923}$ $\underset{(-1.674,0.878)}{-0.398}$
BpAltStarsB [$\lambda=5$] $\underset{(-0.493,1.597)}{0.552}$ $\mathbf{\underset{(0.319,1.706)}{1.013}}$
BpAltKCyclesB [$\lambda=2$] $\underset{(-0.322,0.332)}{0.005}$
BpFourCyclesNodePowerA [$\alpha=1/5$] $\underset{(-4.443,1.662)}{-1.391}$
BpFourCyclesNodePowerB [$\alpha=1/5$] $\mathbf{\underset{(1.670,11.401)}{6.535}}$
The positive and statistically significant BipartiteAltStarsB parameter indicates a tendency for centralisation on plant nodes, that is, plants that attract many pollinators.
The fit of the two models to the cycle length distribution of the network is shown in Fig. 7(a); Model 2 (right) has a slightly better fit to the distribution (and in particular four-cycles) than Model 1 (left). Fig. D1 shows more extensive goodness-of-fit plots for the two models. Model 2 has a better fit to all of the statistics included in the goodness-of-fit plots: degree distributions, giant component size, geodesic distance distribution, dyadwise shared partner distribution, and the bipartite clustering coefficients, as well as the cycle length distribution.

### v3

@@ 7 Simulation experiments
Figure 8: Effect of varying the BipartiteAltKCyclesA parameter (left) and BipartiteFourCyclesNodePowerA parameter (right) on (a) the Edge statistic, (b) the statistic corresponding to the parameter itself and (c) the FourCycles statistic. Each graph plots the mean value of the statistic (over 100 simulations) for three different values of the $\lambda$ or $\alpha$ parameter for BipartiteAltKCyclesA and BipartiteFourCyclesNodePowerA, respectively
8 Empirical example
Here we demonstrate this implementation by using the the statnet ergm package [51] with the new user terms to estimate a model for the Davis “Southern Women” network [91], obtained via the latentnet R package [92, 93]. The network represents the participation of 18 women (first mode) in 14 social events (second mode). This is a well-known affiliation network [94], having been used as an example by Breiger [1] and many papers since, including two [21, 95] in the literature survey (Table LABEL:tab:literature). Although the original publication [91] contains information on the event times, with some exceptions [96, 94] this is not usually used, and we do not use this information here.
The b1np4c and b2np4c model terms take the $\alpha$ ($0<\alpha\leq 1$) value as a parameter, with a default value of $\alpha=0.5$ if omitted. For example, to estimate a model with the b2np4c (BipartiteFourCyclesNodePowerB) term with $\alpha=1/5$ (Model 4 in Table 5):
davis_model4 <- ergm(davis ~ edges + gwb1degree(1, TRUE) +
                             gwb2degree(1, TRUE) + b2np4c(1/5),
Table 5: ERGM parameter estimates for the Southern Women network, estimated with statnet. The table was generated directly from the statnet models with the texreg R package [97].
Model 1 Model 2 Model 3 Model 4
edges $-2.07^{***}$ $-0.20$ $0.47$ $-5.90^{***}$
$(0.34)$ $(0.24)$ $(0.31)$ $(0.58)$
b1star2 $0.07$
$(0.07)$
b2star2 $0.18^{***}$
$(0.04)$
gwb1deg.fixed.1 $-0.83$ $-7.04^{***}$ $10.60^{***}$
$(1.03)$ $(1.51)$ $(1.80)$
gwb2deg.fixed.1 $-2.26^{**}$ $7.76^{**}$ $-6.95^{***}$
$(0.85)$ $(2.48)$ $(1.57)$
gwb1dsp.fixed.0.5 $0.45^{***}$
$(0.12)$
gwb2dsp.fixed.0.5 $-1.33^{***}$
$(0.30)$
b2np4c.fixed.0.2 $17.30^{***}$
$(1.86)$
AIC $319.34$ $328.85$ $308.77$ $285.41$
BIC $329.93$ $339.44$ $326.41$ $299.53$
Log Likelihood $-156.67$ $-161.43$ $-149.38$ $-138.71$
${}^{***}p<0.001$; ${}^{**}p<0.01$; ${}^{*}p<0.05$
Table 5 shows four models for the Southern Women network, estimated with the stochastic approximation algorithm [27]. Model 1, with only the edges and two-star terms for each mode, is the same as [21, Model (8.3)]. Model 2, using the geometrically weighted degree terms rather than two-stars, is similar to [21, Model (8.5)], but using the statnet gwb1degree and gwb2degree terms rather than the BPNet alternating $k$-star terms $\mathrm{K\text{-}S_{P}}$ and $\mathrm{K\text{-}S_{A}}$. Note the reversal of interpretation of signs between $\mathrm{K\text{-}S_{P}}$/$\mathrm{K\text{-}S_{A}}$ and gwb1degree/gwb2degree [98, 99]. Model 3, adding the geometrically weighted dyadwise shared partner terms, is similar to [21, Model (8.6)]. Model 4 uses the new b2np4c term rather than the geometrically weighted dyadwise shared partner terms gwb1dsp and gwb2dsp (models with b1np4c did not converge).
Cycle length distribution goodness-of-fit plots for the four models are shown in Fig. 9, and statnet goodness-of-fit plots in Fig. B1 (Appendix B). Note that all four models fit acceptably well on all the statistics included in the goodness-of-fit tests (degree distributions for each mode, dyadwise shared partners, and geodesic distance distribution), as well as the cycle length distributions. In particular, the fit to four-cycle counts is good for all models, including Model 1 and Model 2, which do not contain any terms to model four-cycles. It therefore appears that Model 1 is the most parsimonious explanation of this data, just as discussed in Wang et al. [21, pp. 22]. This model has a positive and statistically significant event two-star parameter (b2star2), indicating “greater discrepancies in the popularity of events than expected in a random network” [21, pp. 22], taking into account the density (edges) and actor two-star (b1star2) effects.